OpenStudy (anonymous):

simplify: log9 (1/81)^5

OpenStudy (anonymous):

I'm stuck because of the fraction...

OpenStudy (anonymous):

\[\log_9\left(\frac{1}{81}\right)^5=\\\log_9\left(9^{-2}\right)^5=\\\log_99^{-10}=\\-10\]